# Assessing the Adversarial Security of Practical Perceptual Hashing Algorithms

### v1

@@ Abstract
Perceptual hashing algorithms (PHAs) are utilized extensively for identifying illegal online content. Given their crucial role in sensitive applications, understanding their security strengths and weaknesses is critical. This paper compares three major PHAs deployed widely in practice: PhotoDNA, PDQ, and NeuralHash, and assesses their robustness against three typical attacks: normal image editing attacks, malicious adversarial attacks, and hash inversion attacks. Contrary to prevailing studies, this paper reveals that these PHAs exhibit resilience to black-box adversarial attacks when realistic constraints regarding the distortion and query budget are applied, attributed to the unique property of random hash variations. Moreover, this paper illustrates that original images can be reconstructed from the hash bits, raising significant privacy concerns. By comprehensively exposing their security vulnerabilities, this paper contributes to the ongoing efforts aimed at enhancing the security of PHAs for effective deployment.
Index Terms:
Perceptual hashing, security, privacy, blackbox adversarial attack, adversarial machine learning
I Introduction
@@
Perceptual hashing, also known as robust hashing, is a content fingerprinting technique that generates similar binary hashes for perceptually similar contents and different hashes for different contents [1]. Perceptual hashing is useful in various applications such as content moderation, authentication, and detection. Numerous perceptual hash algorithms (PHAs) have been developed or deployed for practical applications [2]. An early example is the algorithm developed by Google and used on YouTube to identify copyrighted material [3]. Other examples include those used in online image search engines of Google, Microsoft Bing, etc [4].
@@
PHAs have gained critical importance, particularly in responding to the reporting of illicit content. For instance, in cases involving Child Sexual Abuse Media (CSAM), US media service providers are legally obliged to examine suspected illegal content and report it to a “Cyber Tip Line” operated by the National Center for Missing and Exploited Children (NCMEC). Microsoft’s PhotoDNA and Meta’s PDQ have been developed for this purpose [5][6][7].
@@
With a growing emphasis on user privacy, technologies like end-to-end encryption have become prevalent. This shift has led to the deployment of PHAs on the client side before data encryption [8]. For example, Apple has proposed to deploy NeuralHash on all its iOS devices [9]. However, deploying PHAs on the client side to scan client data has raised significant concerns regarding client data privacy, prompting widespread criticism about the robustness and privacy-preserving capabilities of the PHAs [10]. Especially, recent studies have almost unanimously criticized PHAs for their vulnerability to adversarial attacks and the potential leakage of information through hash bits [4][11][12].
Given their critical role in sensitive applications, it is imperative to thoroughly understand the security of PHAs. However, even for PHAs that have been widely deployed in the real world, their security remains insufficiently evaluated. Particularly concerning are new PHAs that are hastily deployed without adequate scrutiny. To address this gap, this paper systematically evaluates the adversarial robustness and privacy-preserving capabilities of three widely used PHAs: PhotoDNA, PDQ, and NeuralHash. To support this study, two new efficient attack algorithms are developed.
@@
  • A systematic comparison of the security and privacy-preserving capability of three major PHAs against three types of attacks: normal image editing attacks, malicious adversarial attacks, and hash inversion attacks, with the development of a query-efficient adversarial attack algorithm and a data-efficient hash inversion network.
  • Demonstrating the robustness of PHAs against malicious adversarial attacks when realistic constraints over distortion and query budget are imposed, contrary to previous studies that neglected such constraints and thus criticized PHAs as being not robust. The robustness is shown to be attributed to the unique property of random hash variations.
  • Revealing that the hash bits of all three PHAs leak significant information about the content of the original images, allowing for reliable reconstruction of the original image if sufficient context information is available.
@@
The paper is organized as follows: Section II provides a literature review, Section III formulates the attacks and explains the robustness, Section IV presents the experiment results, and Section V concludes the paper. The source code of the experiments, including the binary executable of the PHAs, can be accessed at https://github.com/neddamj/nnhash.
II Literature Review
II-A Perceptual Hashing Algorithms
Among the vast amount of PHAs [2], this paper considers the following three major ones that have been deployed and used widely in practice:
@@ II-A Perceptual Hashing Algorithms
PhotoDNA, released by Microsoft in 2009, has been used by over 70 companies including Cloudflare and Dropbox [6][13][14]. To create a hash, PhotoDNA breaks down an image into small pieces, analyzes each piece to extract its unique feature, and converts the features into hash bits. PhotoDNA generates a $144$-byte or $1152$-bit hash value for each input image.
PDQ, which stands for Perceptual algorithm utilising a Discrete Cosine Transform and outputting (amongst others) a Quality metric, was released by Meta in 2019 [7]. Inspired by pHash [15], PDQ uses the discrete cosine transform (DCT) to extract image features and outputs both a 256-bit hash value and a quality factor.
NeuralHash was released by Apple in 2021 with the intent to identify CSAM [9]. An image is first passed through a deep neural network (DNN) to produce a feature embedding. The embedding is then converted to a $96$-bit hash. The DNN is trained contrastively to ensure that the cosine similarity between perceptually similar images is high while the similarity of perceptually different images is low.
II-B Adversarial Machine Learning
The security of PHAs is commonly studied using the terminologies and algorithms of adversarial machine learning [16]. Specifically, Adversarial robustness measures the resilience of the machine learning models to adversarial attacks. Adversarial attacks can be classified as untargeted attacks and targeted attacks. In an untargeted attack, the objective is to slightly alter the input to prompt the model to produce a completely different output. Conversely, a targeted attack aims to slightly modify the input to elicit a specifically designed output.
According to the knowledge that the adversary has about the machine learning model, attacks can be categorized to whitebox attacks, where the adversary possesses complete knowledge of the model [16][17][18], and blackbox attacks, where the adversary only has access to the input and output of the model [19][20][21][22]. In black-box attacks, if the model’s outputs are continuous values like logits or probabilities, the attack is referred to as soft-label attack [19][20][21]. Otherwise, if the model’s outputs are solely hard-decision labels, the attack is known as a hard-label attack [22].
II-C Adversarial Robustness of PHAs
The adversarial robustness of conventional PHAs like pHash [2] was investigated in [23] under the whitebox setting, whereas the robustness under the blackbox setting was explored in [4]. Many of these studies suggested that PHAs were not robust to adversarial attacks.
The adversarial robustness of PhotoDNA, PDQ, and NeuralHash has not been adequately addressed, partly because they are new and partly because their source code is not publicly available. Regarding their robustness to normal image editing, [24] examined PhotoDNA, while [11] investigated PDQ and NeuralHash. Both studies demonstrated that PHAs were robust to certain edits like compression but vulnerable to others like cropping.
As far as we know, [25] was the first to study these PHAs’ robustness to adversarial attacks. It studied NeuralHash under the whitebox setting only, and showed that NeuralHash was not robust to whitebox attacks. [26] worked on PDQ and used the linearity of DCT to get adversarial images in a white-box setting. [27] generated adversarial images through linear interpolation between two given images.
Regarding blackbox attacks, [26] studied PDQ and claimed that it was not robust to untargeted blackbox attacks. [28] investigated PhotoDNA and PDQ, concluding that they were not robust to either targeted or untargeted attacks. These two studies closely resemble the current paper, yet this work differs from them in two key aspects. Firstly, previous works computed attack success rates without incorporating limitations on the adversary’s tolerance to image quality degradation or query budget. Such constraints are crucial in real-world scenarios. It is trivial that unconstrained blackbox attacks will always be successful, particularly in the untargeted setting. Secondly, none of the prior works conducted a comparative study of the three PHAs, nor did they investigate the challenge of hash inversion.
II-D Information Leakage from Hashes
Regarding privacy preservation, upon its release, Microsoft asserted that the hashes generated by PhotoDNA were irreversible and incapable of reconstructing the original images [29]. However, in a blog post [12], it was demonstrated that PhotoDNA hashes leak sufficient information about the original image to allow images to be synthesized from the hashes. To our knowledge, [12] is the sole study on reconstructing the original image from the hash, albeit focusing solely on PhotoDNA that has long hashes. It is unclear if PHAs with short hash suffer from this problem.
Another approach to investigate information leakage is to train a classifier to classify the image objects from the hash. [25][27] illustrated that NeuralHash hashes indeed revealed information about their source images by enabling the identification of the object’s class. Interestingly, these findings contradict those of [30] which examined PhotoDNA hashes and claimed that PhotoDNA was resistant to such classification attacks.
III Attack Algorithms
@@
Consider a perceptual hash system where the service provider scans the client’s unencrypted content, extracts a hash for each image using a PHA, compares the hash with a database of known illicit hashes, and raises an alarm or forwards the unencrypted content for further analysis in case of hash match [8]. In this scenario, we consider two adversaries: the client adversary seeking to evade hash matching and the service provider (or some third party) adversary seeking to reconstruct the client’s original images from collected hashes.
We evaluate the security of PhotoDNA, PDQ, and NeuralHash against three types of attacks: benign image editing attacks and malicious adversarial blackbox attacks employed by the client adversary, and hash inversion attacks employed by the service provider adversary. Considering the client adversary, we assess the security of PHAs by measuring the adversary’s attack success rate, where a higher success rate indicates weaker security or robustness. Considering the service provider adversary, security is assessed through image reconstruction quality, where higher quality implies weaker security or privacy.
III-A Image Editing Attacks
While prior works such as [11] and [24] have investigated image editing attacks on PHAs, a new evaluation is still necessary due to ongoing algorithm updates. Notably, NeuralHash has received at least one update since its first release. This study addresses the gap in knowledge by evaluating the security of the updated algorithms and, more importantly, by providing the first comparison of the three PHAs.
III-B Adversarial Blackbox attacks
@@
III-B1 A Query-efficient Blackbox Attack Algorithm
@@
This work focuses solely on blackbox, not whitebox, adversarial attacks, reflecting the practical scenario where PHA implementations are proprietary. For some PHAs, details regarding architecture and model weights are not publicly available due to their commercial nature. For instance, Microsoft enforces non-disclosure agreements for PhotoDNA users, and only an emulation is accessible. Similarly, while Ygvar et al. (2021) released weights for the first-generation NeuralHash, Apple has since updated the algorithm, keeping the new model confidential.
Existing blackbox attack algorithms, designed to operate on logits/probabilities or singular class decisions, cannot be directly applied to PHAs because PHAs output a limited number of hash bits, which are neither logits nor class labels. Instead, we propose a joint soft-label hard-label attack (JSHA) algorithm, where a soft-label attack step is applied first and its output is passed as input to a hard-label attack step. The objective of the soft-label attack step is to modify the image to change the hash without considering the image quality constraint, whereas the objective of the hard-label attack step is to reduce adversarial image quality degradation.
@@
Consider the original image ${\bf I}_{0}$ with hash value ${\bf h}_{0}={\cal H}({\bf I}_{0})$. The objective of the client adversary is to generate an adversarial image ${\bf I}_{\rm adv}$ with new hash ${\bf h}_{\rm adv}={\cal H}({\bf I}_{\rm adv})$. In an untargeted attack, ${\bf h}_{0}$ and ${\bf h}_{\rm adv}$ should be different enough. In a targeted attack, there is a target hash ${\bf h}_{\rm tar}\neq{\bf h}_{0}$, which may be obtained from some target image as ${\bf h}_{\rm tar}={\cal H}({\bf I}_{\rm tar})$, and we require that ${\bf h}_{\rm tar}$ and ${\bf h}_{\rm adv}$ are similar. In both attacks, the distance between ${\bf I}_{0}$ and ${\bf I}_{\rm adv}$ should be small.
To simplify notation, let us use normalized Hamming distance to evaluate the similarity of two $N$-bit hashes as
$D_{h}({\bf h}_{0},{\bf h}_{\rm adv})=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}({\bf h}_{0}(n)={\bf h}_{\rm adv}(n))$ (1)
for two hashes with hash bits ${\bf h}_{0}(n)$ and ${\bf h}_{\rm adv}(n)$, where $\mathbb{I}(x)$ is the indicator function, i.e. $\mathbb{I}(x)=1$ if $x$ is true and 00 otherwise. In addition, we use per-pixel normalized $L_{2}$ norm to evaluate the distance of two images, i.e.
$L_{2}({\bf I}_{0},{\bf I}_{\rm adv})=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\left({\bf I}_{0}(m)-{\bf I}_{\rm adv}(m)\right)^{2}}$ (2)
for two images with pixels ${\bf I}_{0}(m)$ and ${\bf I}_{\rm adv}(m)$. Other measures such as SSIM are also used in our experiments.
@@
The first step of the proposed JSHA algorithm is to apply soft-label blackbox attack to generate an initial adversarial image ${\bf I}_{\rm init}$, with the objective to make the hash ${\bf h}_{\rm init}={\cal H}({\bf I}_{\rm init})$ far away from ${\bf h}_{0}$ for untargeted attacks or close to ${\bf h}_{\rm tar}$ for targeted attacks. This step is formulated as the optimization
$\max_{\delta}\;\;D_{h}\left({\bf h}_{0},{\cal H}({\bf I}_{0}+\delta)\right)$ (3)
in case of untargeted attacks, or
$\min_{\delta}\;\;D_{h}\left({\bf h}_{\rm tar},{\cal H}({\bf I}_{0}+\delta)\right)$ (4)
@@
in case of targeted attacks. The optimization is conducted iteratively. During each iteration, the adversary queries the PHAs with its adversarial input ${\bf I}_{0}+\delta$ to get the hash, and uses the hash to optimize the loss (3) or (4). After reaching the query number threshold or the hash distance threshold, this first step stops with output ${\bf I}_{\rm init}={\bf I}_{0}+\delta$.
@@
[…]
where $p$ is the pre-set hash distance threshold. Starting from ${\bf I}_{\rm adv}={\bf I}_{\rm init}$, the adversary queries the PHAs to get the hash and uses the queried hash to optimize (5) until the pre-set maximum query number or the threshold of $L_{2}({\bf I}_{0},{\bf I}_{\rm adv})$ is reached.
The proposed algorithm is outlined in Algorithm 1. One of its major advantages is query efficiency. A large learning rate can be applied in the first step to get an initial adversarial sample quickly using a small number of queries. In the second step, since we start from ${\bf I}_{\rm init}$ which is only a noisy version of ${\bf I}_{0}$, the convergence is also fast. This makes our algorithm able to generate adversarial samples quickly with a very small number of queries.
Algorithm 1 JSHA: Joint Soft-Label and Hard-Label Attack
1:Initialization: Original input ${\bf I}_{0}$ and its hash ${\bf h}_{0}$; Target hash ${\bf h}_{\rm tar}$ in case of targeted attacks.
2:Step One: Unconstrained soft-label attack
  • The adversary uses queries to optimize (3) or (4)
  • Stop and output ${\bf I}_{\rm init}$ when the pre-set hash distance threshold or query budget is reached.
3:Step Two: Constrained hard-label attack
  • The adversary uses queries to optimize (5)
  • Stop and output ${\bf I}_{\rm adv}$ when the pre-set image distortion threshold or query budget is reached.
@@
Figure 1: (a) Hash inversion attack model. (b) Architecture of residual block.
III-B2 Robustness to Adversarial Blackbox Attacks
Algorithm 1 and most existing blackbox attack algorithms rely on gradient estimation to solve the optimization problem. Drawing on the property of PHAs and our prior work on adversarial robust analysis [31], we have the following proposition.

### v2

# Robustness of Practical Perceptual Hashing Algorithms to Hash-Evasion and Hash-Inversion Attacks

@@ Abstract
Perceptual hashing algorithms (PHAs) are widely used for identifying illegal online content and are thus integral to various sensitive applications. However, due to their hasty deployment in real-world scenarios, their adversarial security has not been thoroughly evaluated. This paper assesses the security of three widely utilized PHAs—PhotoDNA, PDQ, and NeuralHash—against hash-evasion and hash-inversion attacks. Contrary to existing literature, our findings indicate that these PHAs demonstrate significant robustness against such attacks. We provide an explanation for these differing results, highlighting that the inherent robustness is partially due to the random hash variations characteristic of PHAs. Additionally, we propose a defense method that enhances security by intentionally introducing perturbations into the hashes.
1 Introduction
@@
Perceptual hashing, also known as robust hashing, is a content fingerprinting technique that generates similar binary hashes for perceptually similar contents and different hashes for different contents [1]. Perceptual hashing is useful in various applications such as content moderation, authentication, and detection. Numerous perceptual hash algorithms (PHAs) have been developed or deployed for practical applications [2]. Early examples include the algorithm used on YouTube to identify copyrighted material [3] and algorithms used in online image search engines of Google, Microsoft Bing, etc [4].
@@
Recently, PHAs have gained more critical importance, particularly in responding to the requiremnt of reporting illicit content. For instance, in cases involving Child Sexual Abuse Media (CSAM), US media service providers are legally obliged to examine suspected illegal content and report it to a “Cyber Tip Line” operated by the National Center for Missing and Exploited Children (NCMEC). Microsoft’s PhotoDNA and Meta’s PDQ were immediately deployed for this purpose [5][6][7].
@@
With a growing emphasis on user privacy, technologies like end-to-end encryption have become prevalent. This shift has led to the deployment of PHAs on the client side before data encryption [8]. For example, Apple has proposed to deploy NeuralHash on all its iOS devices [9]. However, this client-side deployment has raised significant concerns among the client users about their data privacy, which inadvertently arouses a widespread criticism regarding the security and privacy-preserving capabilities of PHAs [10]. Criticism has been further fueled by recent studies highlighting the vulnerability of PHAs to adversarial attacks and the potential for information leakage through hash bits [4][11][12].
Given the critical role PHAs play in sensitive applications, it is essential to ensure their security. Unfortunately, PHAs are often deployed hastily without sufficient security evaluation. This paper systematically assesses the adversarial robustness of three widely used PHAs—PhotoDNA, PDQ, and NeuralHash—against two types of attacks: hash-evasion attacks, aimed at enabling illicit content to evade detection, and hash-inversion attacks, aimed at reconstructing the original secret content from hashes. To support this evaluation, we introduce two new attack algorithms: a query-efficient black-box adversarial attack algorithm and a data-efficient hash inversion algorithm.
@@
  • This study demonstrates that PHAs exhibit notable robustness against blackbox adversarial evasion attacks when realistic constraints on image distortion, query budget, and hash difference are applied. This finding challenges previous studies that overlooked these constraints, leading to unjustified criticisms of PHAs.
  • This paper confirms the sufficient privacy-preserving capabilities of hash bits, countering claims from some existing studies that suggested hash bits disclose significant information about the content of the original images. The key point is how diverse the images are.
  • This research reveals that the security of PHAs is partially due to the unique property of random hash variations. Based on this observation, a defense method is proposed to enhance security by intentionally randomizing hash bits.
@@
The paper is organized as follows: Section 2 provides a literature review, Section 3 formulates the attack and defense algorithms, Section 4 presents the experiment results, and Section 5 concludes the paper. More detailed experiment data can be found in the Appendix. The source code of the experiments, including ways to access binary executables of the PHAs, can be found at https://github.com/neddamj/nnhash.
2 Literature Review
Among the numerous PHA’s [2], this paper focuses on the following three major ones due to their widespread deployment in practice:
@@ 2 Literature Review
  • PhotoDNA, released by Microsoft in 2009, is used by over 70 companies including Cloudflare and Dropbox [6][13][14]. It generates an 1152-bit hash for each input image;
  • PDQ, launched by Meta in 2019 [7] and heavily inspired by pHash [2][15], produces a 256-bit hash for each input image.
  • NeuralHash, released by Apple in 2021 with the goal of identifying CSAM [9], creates a more concise 96-bit hash for each input image.
The adversarial robustness of PhotoDNA, PDQ, and NeuralHash to either hash-evasion or hash-inversion attacks has not been thoroughly investigated, primarily due to their relative novelty and the absence of publicly available source code. For hash-evasion, [16] was the first to explore the robustness of NeuralHash, revealing its vulnerability to whitebox attacks, assuming the adversary has full access to the PHA. Additionally, [17] examined PDQ and demonstrated how adversarial images could be generated in the whitebox context. Furthermore, [18] utilized linear interpolation between two images to create adversarial examples.
Regarding blackbox hash-evasion attacks, where the adversary only has access to the input and output of the algorithm, [17] examined PDQ. It concluded that it was not robust against untargeted blackbox attacks. Similarly, [19] evaluated both PhotoDNA and PDQ, finding that they lacked robustness against both targeted and untargeted attacks. While these studies are closely related to the current paper, they overlooked the importance of adversaries’ constraints on image distortion, query budget, as well as hash difference levels. By all means, unconstrained blackbox attacks may always be successful, particularly in the untargeted setting.
Regarding hash-inversion attacks or privacy-preserving capabilities of PHAs, Microsoft initially claimed that the hashes generated by PhotoDNA were irreversible and could not be used to reconstruct the original images [20]. However, a blog post [12] challenged this assertion, suggesting that PhotoDNA hashes might contain enough information to synthesize images from the hashes. To the best of our knowledge, [12] is the only work addressing the reconstruction of original images from hashes, and it specifically focuses on PhotoDNA, which uses long hashes. Another method for investigating information leakage involves training classifiers to identify image content from hashes. Studies such as [16] and [18] demonstrated that NeuralHash hashes could indeed reveal information about their source images, such as identifying the object’s class. These findings contrast with those of [21], which examined PhotoDNA hashes and concluded that they were resistant to such classification attacks.
3 Attack Algorithms and Defense Method
@@
In a perceptual hash system, a service provider scans a client’s unencrypted content, extracts hashes for each image using a perceptual hashing algorithm (PHA) and compares these hashes with a database of known illicit hashes. If a match is found, an alarm is raised or the unencrypted content is forwarded for further analysis [8]. In this context, we consider two types of adversaries: the client adversary, who aims to evade hash matching, and the service provider (or third-party) adversary, who seeks to reconstruct the client’s original images from collected hashes. This paper evaluates the security of three widely used PHAs—PhotoDNA, PDQ, and NeuralHash—against two specific types of attacks: blackbox hash-evasion attacks, which are employed by the client adversary, and hash-inversion attacks, which are utilized by the service provider adversary. Additionally, we propose a defense mechanism that can be integrated into the perceptual hashing system.
@@
3.1 A Query-Efficient Adversarial Blackbox Attack Algorithm for Hash Evasion
@@
This study focuses exclusively on blackbox adversarial attacks, rather than whitebox attacks, to reflect the practical scenario where PHA implementations are proprietary. In many cases, the details of PHA architectures and model weights are not publicly available due to their commercial nature. For example, Microsoft enforces non-disclosure agreements for PhotoDNA users, providing only an emulation of the algorithm. Similarly, although [22] released weights for the first-generation NeuralHash, Apple has since updated the algorithm, and the details of the latest version remain confidential.
@@
Consider the original image ${\bf I}_{0}$ with the hash value ${\bf h}_{0}={\cal H}({\bf I}_{0})$. The objective of the client adversary is to generate a very similar image ${\bf I}_{\rm adv}$ whose new hash ${\bf h}_{\rm adv}={\cal H}({\bf I}_{\rm adv})$ is sufficiently different from ${\bf h}_{0}$ in untargeted attacks or sufficiently similar to some target hash ${\bf h}_{\rm tar}$ in targeted attacks. We use normalized Hamming distance $D_{h}({\bf h}_{0},{\bf h}_{\rm adv})$ to evaluate the similarity of two hashes, and use per-pixel normalized $L_{2}$ norm $L_{2}({\bf I}_{0},{\bf I}_{\rm adv})$ to evaluate the similarity of two images.
@@
Because the output of PHAs is hash bits, not logits or class decisions, we propose a joint soft-label hard-label attack (JSHA) algorithm that can drastically improve blackbox attack query efficiency. JSHA consists of two steps. The first step is to apply a soft-label blackbox attack to generate an initial adversarial image ${\bf I}_{\rm init}$ so that its hash ${\bf h}_{\rm init}={\cal H}({\bf I}_{\rm init})$ is far away from ${\bf h}_{0}$ for untargeted attacks or close to ${\bf h}_{\rm tar}$ for targeted attacks. This step is formulated as the optimization
[…]
@@
The optimization is conducted iteratively. In each iteration, the adversary queries the PHAs with its adversarial input ${\bf I}_{0}+\delta$ to get the hash, and uses the hash to optimize the loss. After reaching the query number threshold or the hash distance threshold $D_{0}$, this first step stops with output ${\bf I}_{\rm init}={\bf I}_{0}+\delta$.
@@
Starting from ${\bf I}_{\rm adv}={\bf I}_{\rm init}$, the adversary queries the PHAs to get the hash and uses the queried hash to optimize (2) until the query number threshold or the threshold of $L_{2}({\bf I}_{0},{\bf I}_{\rm adv})$ is reached.
One of major advantages of this algorithm is high query efficiency. The first step is an unconstrained optimization and a large learning rate can be applied to get an initial adversarial sample quickly. In the second step, since we start from ${\bf I}_{\rm init}$ which is only a noisy version of ${\bf I}_{0}$, the convergence is fast.
@@
Figure 1: (a) Hash inversion attack model (left: for Celeb Dataset. Right: for STL-10 dataset). (b) Architecture of residual block.
3.2 A Data-Efficient Hash-Inversion Attack Algorithm
GAN (generative adversarial network) was trained in [12] to invert CelebA images. However, the adversary may need more data-efficient hash-inversion attack algorithms because the training data may not be abundant. For this, we has developed new and more data-efficient generative models to reconstruct the original input images from the hashes. We designed a decoder-style deep model, similar to the Fast Style Transfer [23], to convert the $N$-bit hashes to $(64,64)$ images. This model is mainly a combination of residual blocks depicted in Fig. 1(b), with fractionally-strided convolutions/deconvolutions for learned upsampling. The entire architecture is described in Fig. 1(a), where the left network is for the MNIST and CelebA dataset while the right network is used for the STL-10 dataset.
For a given training dataset, we first compute the hashes of each image offline and store the image-hash pairs. We then train the model outlined above to reconstruct the images as closely as possible by minimizing the mean-square-error (MSE) loss between the model output and the image from which the hash was generated.
3.3 A Defense Method based on Random Hash Perturbation
The adversaries rely on gradient estimation to solve the optimization problems. It is well known that hash bits of PHAs change slightly and randomly with respect to slight changes in input images. Based on existing work on random perturbation defense [24][25], we have the following proposition: